OpenStudy (anonymous):

what are the volume formulas for these? 1.cylinder 2.cone 3.sphere

OpenStudy (anonymous):

@surjithayer ?

OpenStudy (anonymous):

are you looking for area volume or what?

OpenStudy (anonymous):

oooooh volume sorry

OpenStudy (anonymous):

\[cylinder\] curved surface =2 pi r h r=radius of the base of cylinder h=height Volume=pi r^2 h

OpenStudy (anonymous):

Ok so you can easily google this but I will tell you

OpenStudy (anonymous):

area of base=pi r^2

OpenStudy (anonymous):

a cone is\[\Pi r^2(\frac{ h }{ 3 })\]

OpenStudy (anonymous):

2.cone curved surface=pi r l r=radius l=slant height of cone Volume\[=\frac{ 1 }{3 }\pi r^2h \]

OpenStudy (anonymous):

\[\frac{ 4 }{ 3 } \Pi r^3\] is a sphere

OpenStudy (anonymous):

sphere surface area=4 pi r^2 \[Volume=\frac{ 4 }{ 3 }\pi r^3\]

OpenStudy (anonymous):

THANKS XD

OpenStudy (anonymous):

Np

OpenStudy (anonymous):

yw